Question 16

Consider the following quantified statement:

∀x ∈ Z [(x2 ≥ 0) ∨ (x2 + 2x – 8>0)].

Which one of the alternatives provides a true statement regarding the given statement or its 

negation?

 

1. The negation ∃x ∈ Z [(x2 < 0) ∨ (x2 + 2x – 8 ≤ 0)] is not true.

2. x = – 3 would be a counterexample to prove that the negation is not true.

3. x = – 6 would be a counterexample to prove that the statement is not true.

4. The negation ∃x ∈ Z [(x2 < 0) ∧ (x2 + 2x – 8 ≤ 0)] is true.


1
Expert's answer
2021-07-22T09:41:32-0400

Consider the following quantified statement "∀x ∈ \Z [(x^2 ≥ 0) ∨ (x^2 + 2x – 8＞0)]." The negation is "\exists x ∈ \Z [(x^2 < 0) \land (x^2 + 2x – 8\le 0)]."


  1. Since "x^2\ge 0" for any "x\in\Z," we conclude that the negation "\exists x ∈ \Z [(x^2 < 0) \land (x^2 + 2x – 8\le 0)]" is not true.
  2. For "x=-3" we have that "(-3)^2=9\ge0" and "(-3)^2+2(-3)-8=-5<0", and hence "[((-3)^2 < 0) \land ((-3)^2 + 2(-3) – 8 ≤ 0)]" is not true. But this is not a counterexample because of in the prefix of the negation there is the existence quantor.
  3. For "x=-6" we have that "(-6)^2=36\ge0" and "(-6)^2+2(-6)-8=16>0", and hence "[((-6)^2 \ge 0) ∨ ((-6)^2 + 2(-6) – 8 > 0)]" is true.
  4. Since "x^2\ge 0" for any "x\in\Z," we conclude that the negation "\exists x ∈ \Z [(x^2 < 0) \land (x^2 + 2x – 8\le 0)]" is false.

Answer: 1